OpenStudy (anonymous):

Given the equation y − 3 = 1 half (x + 6) in point-slope form, identify the equation of the same line in standard form. please tell me if its a. x − 2y = −12 b. y = 1/2x c. y = 1/2x + 6 d. y = 1/2x + 9 please give an answer

OpenStudy (anonymous):

y = mx+c Standard Form..)

OpenStudy (anonymous):

Standard form is ax+by=c Slope-intercept form is y=mx+b You can convert the equation of this line into slope-intercept form to get answer C. y - 3 = 1/2 (x + 6) y = 1/2x + 3 + 3 y = 1/2x + 6

OpenStudy (anonymous):

But thats not standard form

OpenStudy (anonymous):

Only one of the answers is in standard form (A.), and if you convert A. into slope-intercept form, you get the same equation as C. x - 2y = -12 -2y = -x - 12 -2 _______________ y = 1/2x + 6